OpenStudy (anonymous):

Help anyone 5 questions only Answers just need correction if needed

OpenStudy (anonymous):

1.b 2.a 3.d 4.c 5.c

OpenStudy (anonymous):

@jim_thompson5910 anyone?

OpenStudy (lasttccasey):

1. A 2. B 3. C 4. D 5. B

OpenStudy (lasttccasey):

Do you know how to use the quadratic formula?

OpenStudy (anonymous):

yes i get confused sometimes but i know how to use it its -B+-sqrtb^2-4ac/2a right?

OpenStudy (lasttccasey):

Yes\[\frac{ -B \pm \sqrt{B^2-4AC} }{ 2A }\]

OpenStudy (anonymous):

I have 5 more with answers can you help me ?

OpenStudy (lasttccasey):

Sure but I want you to know how to do this. Do you want me to walk through the 1st one with you real quick?

OpenStudy (anonymous):

sure lets walk threw the 1st of the other ones k

OpenStudy (anonymous):

1.not sure whats the answer 2.b or a 3 c 4 a 5.c or b

OpenStudy (lasttccasey):

This website goes through the process fairly well for these problems: http://hotmath.com/hotmath_help/topics/finding-the-equation-of-a-parabola-given-focus-and-directrix.html Its easier to follow this for the parabolic equations that trying to explain it through text. I'll see if I can get the answers real quick though.

OpenStudy (anonymous):

k ill take notes on the subject while you do that thanks by the way :)

OpenStudy (lasttccasey):

Okay and its not a problem. 1. C 2. D 3. C 4. B 5. D

OpenStudy (anonymous):

thanks for the help :)

OpenStudy (lasttccasey):

Basically (x-a)^2+(y-b)^2=(y-d)^2, where (a,b) are the focus and y=d is the directrix. Hope that helps!

OpenStudy (anonymous):

yes it does :)

OpenStudy (lasttccasey):

You're welcome :)